A solution is prepared from 20mL of 0.10M HNO3 and 30mL of 0.05M H2SO4. If the H2SO4 completely ionises, the pH of the solution would be:
A. +1 B. +3 C. –log 0.02 D. –log 0.07
1
Expert's answer
2020-08-17T13:34:55-0400
"n_1(H^+) = c(HNO_3)V_1 = 0.10M*0.02L = 0.002mol"

"n_2(H^+)=2c(H_2SO_4)V_2 = 2*0.05M*0.03L = 0.003mol"

"n_{total} = 0.002mol+0.003mol = 0.005mol"

"V_{total} = 0.02L+0.03L = 0.05L"

"c(H^+) = n\/V = 0.005mol\/0.05L = 0.10M"

"pH = -logc(H^+) = +1"

Answer: A +1